# EDTformer: An Efficient Decoder Transformer for Visual Place Recognition

### v1

@@ IV-D Ablation Studies
Ablation on the rank of LoPA. In this subsection, we further conduct ablation studies about the effect of the rank in LoPA. We consistently use the EDTformer for feature aggregation. Table X shows the results of setting different ranks. We set “w/o LoPA” as baseline. Setting the rank to 4 gets the best overall performance, but we can still achieve good results with rank even set to 2. These results show that we can achieve an outstanding performance without introducing a large number of parameters and consuming substantial memory during training. It is worth mentioning that the backbone DINOv2 is completely frozen, which indicates that the intermediate features produced by each block of DINOv2 remain consistent. The final output solely depends on how these intermediate activations are processed, which enlightens our future work to design a better parallel network to refine them. Meanwhile, due to only introducing a few parameters and memory usage, it can facilitate the application of vision foundation models under resource-limited conditions.
@@ IV-E Qualitative experiments
Fig. 5 presents the qualitative experimental results in various challenging environments, including viewpoint variants, drastic condition changes, severe occlusions, image domain variations, etc. In the vast majority of cases, our method successfully gets similar and correct place images from the same location as the query images, while other methods tend to struggle with obtaining the correct images within the predefined threshold (25m), which demonstrates that our method is highly robust to environment changes and less prone to perceptual aliasing. For instance, in the seventh and eighth examples, the query images are captured at night. Only a small region of the image is clearly visible. However, our method still gets the correct results. Additionally, it is worth mentioning that all methods fail to obtain the correct places in the last two examples, where multiple challenges (e.g., viewpoint changes, domain variations, severe occlusions, dynamic objects and perceptual aliasing) arise simultaneously. This motivates us to further improve our method in future work to get higher accuracy and tackle more challenging scenarios.
@@ IV-F Interpretability analysis
The experimental results presented in Table II and III have fully demonstrated the superior performance of our method. In this section, we further explore the underlying reasons. To this end, we visualize the attention weights between the input image and the learnable queries within the EDTformer. We randomly highlight two learnable queries from the outputs of the second decoder block to understand their unique learning characteristics. The results are shown in Fig. 6. Intuitively, each query has a different focus, but they all share a common characteristic. Observing horizontally, we can find that each query basically focuses on invariant yet discriminative regions (e.g., buildings and trees) while ignoring useless, dynamic and easily changing objects (e.g., sky, vehicles and pedestrians). In other words, our EDTformer successfully decodes crucial features for the VPR task into these learnable queries to achieve feature aggregation. As a result, with all information within the learnable queries combined, they can consistently concentrate on the majority of discriminative regions in the place images, which is why our method can effectively address various challenging issues in VPR.
@@ V Conclusions
In this paper, we revisit the transformer decoder and propose a novel feature aggregation method EDTformer, a simple, efficient and powerful decoder architecture which can directly generate a robust and discriminative global representation for the place image. To be specific, EDTformer takes deep features extracted from the backbone as the keys and values, as well as a set of independent learnable queries as the queries. Then it utilizes the attention mechanism to decode and aggregate the effective features into the learnable queries, resulting in the final global representation. To further enhance the robustness, we introduce the vision foundation model DINOv2 as the backbone and design a Low-rank Parallel Adaptation method, which can adapt DINOv2 to the VPR task in a memory- and parameter-efficient way to improve the performance of the whole model. The extensive experiments show that our method effectively addresses various challenging problems in VPR and outperforms other SOTA methods on multiple benchmark datasets by a large margin.

### v2

@@ IV-D Ablation Studies
Effect of the rank of LoPA. In this subsection, we further conduct ablation studies about the effect of the rank in LoPA. We consistently use the EDTformer for feature aggregation. Table X shows the results of setting different ranks. We set “w/o LoPA” as baseline. Setting the rank to 4 gets the best overall performance, but we can still achieve good results with rank even set to 2. These results show that we can achieve an outstanding performance without introducing a large number of parameters and consuming substantial memory during training. It is worth mentioning that the backbone DINOv2 is completely frozen, which indicates that the intermediate features produced by each block of DINOv2 remain consistent. The final output solely depends on how these intermediate activations are processed, which enlightens our future work to design a better parallel network to refine them. Meanwhile, due to only introducing a few parameters and memory usage, it can facilitate the application of vision foundation models under resource-limited conditions.
Effect of LoPA on other VPR methods. To further demonstrate the universality of our proposed LoPA, we conduct ablation studies about the effect of LoPA on some other common VPR methods, including GeM, NetVLAD, and SALAD. The results are reported in Table XI. Only using the DINOv2-base backbone without LoPA to adapt it for the VPR task leads to poor performance. However, with the help of LoPA, all these VPR methods can achieve a significant improvement. For example, GeM with LoPA achieves nearly 3$\times$ higher R@1 on Nordland compared to its baseline. These results indicate that our proposed LoPA is not tailored for our EDTformer, but also is applicable to some other VPR methods.
@@ IV-E Qualitative Experiments
Fig. 5 presents the qualitative experimental results in various challenging environments, including viewpoint variants, drastic condition changes, severe occlusions, image domain variations, etc. In the vast majority of cases, our method successfully gets similar and correct place images from the same location as the query images, while other methods tend to struggle with obtaining the correct images within the predefined threshold (25m), which demonstrates that our method is highly robust to environment and viewpoint changes, as well as less prone to perceptual aliasing. For instance, in the seventh and eighth examples, the query images are captured at night. Only a small region of the image is clearly visible. However, our method still gets the correct results. Additionally, it is worth mentioning that all methods fail to obtain the correct places in the last two examples, where multiple challenges (e.g., viewpoint changes, domain variations, severe occlusions, dynamic objects and perceptual aliasing) arise simultaneously. This motivates us to further improve our method in future work to get higher accuracy and tackle more challenging scenarios.
@@ IV-F Interpretability Analysis
The experimental results presented in Table II and III have fully demonstrated the superior performance of our method. In this subsection, we further explore the underlying reasons. To this end, we visualize the attention weights between the input image and learnable queries within our EDTformer to understand their unique aggregating characteristics. The results are shown in Fig. 6. Each single learnable query has a different focus, but they all share a common characteristic. Observing horizontally (especially the first two examples), we can find that each learnable query basically focuses on short-term invariant yet discriminative regions (e.g., buildings, trees, and some special signs on the ground) while ignoring dynamic and easily changing objects (e.g., sky, vehicles, and pedestrians) that are useless for the VPR task [70, 4]. Moreover, when facing various challenges (e.g., illumination and seasonal changes, viewpoint changes, etc) as shown in the last four examples, our method can also tackle them well. These results indicate that our EDTformer successfully decodes crucial features for the VPR task into these learnable queries to achieve feature aggregation. With all information within the learnable queries combined, they can consistently concentrate on the majority of discriminative regions in the place images, which is why our method can achieve a great recognition performance and effectively address most challenges in VPR.
@@ V Conclusions
In this paper, we revisited the transformer decoder and proposed a novel feature aggregation method EDTformer for VPR, which is simple, efficient and powerful. The EDTformer, primarily consisting of our simplified decoder blocks, can fully decode and aggregate the effective information within deep features extracted from the backbone into a set of learnable queries, thus producing a robust and discriminative global representation for the place image. Moreover, we designed a Low-rank Parallel Adaptation method, which can adapt the foundation model DINOv2 for the VPR task in a memory- and parameter-efficient way, thereby providing more powerful features for our EDTformer and enhancing the robustness of the entire model. The extensive experiments show that our approach surpasses other SOTA methods on multiple benchmark datasets by a large margin with less training memory usage and effectively tackles various challenges in VPR.